# Long-Tailed 3D Detection via 2D Late Fusion

### v3

@@ 4.1 Comparison to the State-of-the-Art Methods
Table 2: Fusing uni-modal detections in the 3D BEV vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchical loss [5]) with 2D RGB and 3D RGB detections in both the 3D BEV and 2D image plane. We match and filter detections in the 3D BEV using center distance as proposed by [5], and match and filter detections in the 2D image plane using IoU. Following [9], we inflate 2D detections to the 3D BEV using LiDAR points within the box frustum. We project 3D detections to the 2D plane using provided sensor extrinsics. We find that matching 3D RGB detections in the 3D BEV and in the 2D image plane yields similar results. Unsurprisingly, inflating 2D RGB detections for matching in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV. In contrast, filtering LiDAR-based detections using 2D detections in the 2D image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP. This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors. We further verify this in Table 9.
 Method Fusion in 3D BEV Fusion on 2D Image Plane
@@
4.2 Ablation Study
In this section, we design a set of experiments to study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image and 3D BEV plane (cf. Table 2). Further, we examine the impact of using additional data and study different fusion strategies (cf. Table 3).
How do we incorporate RGB information? Although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs using an RGB-based detector by leveraging two insights: LiDAR-based 3D detections are accurate w.r.t 3D localization and yield high recall (though classification is poor), and RGB-based detections are accurate w.r.t recognition (though 3D localization is poor). We incorporate RGB information by matching and filtering 3D LiDAR detections with RGB-based detections. We evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO Det.) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer) in Table 2. We also analyze the impact of matching LiDAR detections with 2D RGB detections in the 3D BEV and 2D image plane. As shown in Table 2, matching LiDAR detections with 2D RGB detections in the 2D image plane (bottom right) performs best.
Table 3: Ablation on late-fusion. Our analysis confirms that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration prior and probabilistic fusion improve performance. Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO. In addition, training with external data further improves rare class detection by 4.2%.
@@ 4.2 Ablation Study
       CenterPoint w/ Hierarchy [5]        40.4        77.1        45.1        4.3
        + 2D Img. Filtering w/ DINO Det.        +7.5        47.9        77.1        55.8        14.4
@@
How do we match detections from uni-modal detectors? To match and filter LiDAR and RGB detections in the 3D BEV, we follow the procedure introduced by [5]. For each RGB-based detection, we keep LiDAR-based detections within a radius of $m$ meters and remove all the others that are not close to any RGB-based detections. This works well for 3D RGB detections. Late fusion with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only model by 2% averaged over all classes (cf. Table 2). Unsurprisingly, we find that matching inflated 2D RGB detections in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV, notably achieving marginally lower accuracy than the LiDAR-only baseline.
@@
Table 2 shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10% mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that 2D detectors achieve better recognition performance than 3D RGB detectors (cf. Table 9).
How do we fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7% averaged over all classes, notably improving performance for classes with a medium number of examples (cf. Table 3). For probabilistic ensembling, we use Bayesian fusion to reason about the final score of matched detections. Concretely, if two matched detections fire in the same place, the fused score should be higher than the individual scores because there is twice the evidence of an object at that particular spatial location. As shown in Table 3, this further improves accuracy by 0.9 mAP.
@@ 5 Conclusion
In this paper, we present a detailed exploration of late-fusion for long-tailed 3D detection. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration and probabilistic fusion improve performance. Our simple late-fusion approach achieves state-of-the-art performance, improving over prior work by 5.9% mAP.
Appendix A Implementation Details.
@@
We include a pipeline diagram of our late-fusion approach above for clarity. We retrain several existing LiDAR-only, RGB-only, and multi-modal detectors on all 18 classes in the nuScenes LT3D setup. We employ standard augmentation techniques (such as copy-paste augmentation) and adopt the sampling ratios defined by [5] when training all detectors. Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We use the open-source implementations from [5] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper. We train our model with 8 RTX 3090 GPUs. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [68] and COCO [50] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively.
@@
Table 4: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2% overall and CMT-L by 15.9%. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
@@
          CenterPoint w/ Hier. + DINO Det.           51.4           77.9           59.4           20.0
@@
Appendix B Transformer-Based LiDAR Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [5]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [6] and CMT-L [66] in Table 4.
First, we note that BEVFusion-L performs 2.1% better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%. However, BEVFusion-L performs 4.6% worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging. Surprisingly, CMT-L performs worse than CenterPoint on all classes. We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2% improvement overall, and a 8.5% AP improvement for rare classes. This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy. In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3% overall, and performs 5.2% worse than our late-fusion approach. Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9% improvement overall, with a 19.1% improvement for rare classes. Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
Table 5: NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
FCOS3D BEVFormer CP TransFusion DeepInteraction CMT CP+FCOS3D CP+DINO
[7] [21] [8] [31] [65] [66] [5] (Ours)
mAP 20.9 27.3 39.2 39.8 43.7 44.4 43.6 51.4
NDS 30.4 38.8 54.9 53.9 54.4 55.9 56.7 60.4
@@ Appendix B Transformer-Based LiDAR Detectors
Table 6: Comparison of inference timing (in ms). Our late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from Table 1 for reference, demonstrating that our simple late-fusion approach resoundingly outperform sophisticated state-of-the-art methods (e.g., Deep Interaction and TransFusion) with significantly less inference time!
FCOS3D BEVFormer CP TransFusion DeepInteraction CP+FCOS3D CP + DINO
[7] [21] [8] [31] [65] [5] (Ours)
Time (ms) 89 327 323 367 590 323 323
mAP 20.9 27.3 39.2 39.8 43.7 43.6 51.4
Appendix C AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table 5. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.

### v4

@@ 5.2.1 Analysis on Training with Semantic Hierarchy
TABLE III: Fusing uni-modal detections in 3D vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchy loss) with 2D RGB and 3D RGB detections in 3D versus on the image plane. We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF). Following [70], we inflate 2D detections to 3D using LiDAR points within the box frustum. We project 3D detections to the image plane using provided sensor extrinsics. Results show that matching 3D RGB detections in 3D and on the image plane yields similar results. Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D. In contrast, filtering LiDAR-based detections using 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP. This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors. We further verify this in the supplement.
 Method Fusion in 3D (MMF) Fusion in 2D (MMLF)
@@
TABLE IV: Progressively adding training components improves LT3D performance, confirming that 2D RGB detectors are better suited for multi-modal late-fusion, matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and score calibration prior and probabilistic fusion improves performance. Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO. In addition, training with external data further improves rare class accuracy by 4.2 mAP.
@@
CenterPoint [12] 39.2 76.4 43.1 3.5
 + Hierarchy Loss +1.2 40.4 77.1 45.1 4.3
 + MMF w/ DINO +7.5 47.9 77.1 55.8 14.4
@@
Figure 10: Both our MMLF method (columns 1-3) and TransFusion [34] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult.
5.2.2 Analysis on Multi-Modal Late-Fusion
We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table III. We further examine the impact of using additional data and study different fusion strategies in Table IV.
How do we incorporate RGB information? Table II demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D. We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table III. Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
@@ 5.2.2 Analysis on Multi-Modal Late-Fusion
How do we match detections from uni-modal detectors? MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections. This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table III. Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, notably achieving marginally lower accuracy than the LiDAR-only baseline.
@@
Table III shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10 mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that MMLF performs better with 2D RGB detectors than 3D RGB detectors.
How do we fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7 mAP averaged over all classes (Table IV). Once calibrated, we probabilistically fuse matched detections to generate the final set of detections. Table IV shows that probabilistic fusion improves performance by 0.9 mAP.
TABLE V: Diagnosis using the mAPH metric on selected classes. We analyze CenterPoint [12] and CMT [10], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [11]). Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
CenterPoint [12] LCA=0 82.4 81.2 49.4 19.7 33.6 48.9 0.1 14.2 0.1 21.7
(original) LCA=1 83.9 82.0 58.7 20.5 35.2 50.5 0.1 18.3 0.1 22.0
LCA=2 84.0 82.4 58.8 20.7 36.4 51.0 0.1 19.5 0.1 22.6
CenterPoint LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
(Group-Free) LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
LCA=2 89.1 87.5 71.7 26.8 51.1 65.2 0.1 30.5 4.8 33.4
CenterPoint LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
(Group-Free) LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
+ Hierarchy Loss LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
(Group-Free) LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
+ Hier. + MMLF LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
  CMT [10] LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
(LiDAR-only) LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
+ MMLF LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
5.2.3 Analysis of Misclassification
We evaluate CenterPoint [12] and CMT [10] using the hierarchical mAP metric (mAPH), which considers semantic relationships across classes to award partial credit. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child). Table V lists the results of applying our developed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO. Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
5.2.4 Analysis of Failure Cases
We visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [34], an end-to-end trained multi-modal detector. We find that our MMLF fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects. See Fig. 10 for detailed analysis.
Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm.
5.3 Limitations and Future Work
Our work does not directly study how addressing LT3D affects downstream perception tasks. Future work should address this limitation. Moreover, as shown in Fig. 11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. We find that 2D detection accuracy is a strong proxy for final 3D LT3D performance. Recent works in large-scale vision language models [77, 78, 79, 80, 81] show promising zero-shot results in detecting rare classes. Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further. In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras). Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another. If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections. In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
@@ 6 Conclusion
We explore the problem of long-tailed 3D detection (LT3D), detecting objects not only from common classes but also from many rare classes. This problem is motivated by the operational safety of autonomous vehicles (AVs) but has broad robotic applications, e.g., elder-assistive robots [82] that fetch diverse items [83] should address LT3D. To study LT3D, we establish rigorous evaluation protocols that allow for partial credit to better diagnose 3D detectors. We propose several algorithmic innovations to improve LT3D, including a group-free detector head, a hierarchical loss that promotes feature sharing across long-tailed classes, and present a detailed exploration of multi-modal late-fusion for LT3D. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance. Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes!
Acknowledgments
This work was supported by the University of Macau (SRG2023-00044-FST), NSFC (No.62206256), CMU Center for Autonomous Vehicle Research, and NSF GRFP (DGE2140739).
A Implementation Details
@@
We use the open-source implementations from [9] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [84] and COCO [55] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively. We describe important implementation details below.
  • Input. We adopt 10-frame aggregation for LiDAR densification when training LiDAR-based detectors on nuScenes and a 5-frame aggregation on Argoverse 2. We assume that we are provided with ego-vehicle poses for prior frames to align all LiDAR sweeps to the current ego-vehicle pose. Since LiDAR returns are sparse, this densification step is essential for accurate 3D detection. By default, we train the 2D RGB detectors on the 2D bounding boxes derived by projecting 3D annotations to the 2D image plane and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors YOLOV7 and DINO are pre-trained on the ImageNet [84] and COCO [85] datasets.
  • Model Architecture. We adopt the architecture in [18] but make an important modification. The original architecture (for the standard nuScenes benchmark) has six heads designed for ten classes; each head has 64 filters. We first adapted this architecture for LT3D using seven heads designed for 18 classes. We then replace these seven heads with a single head consisting of 512 filters shared by all classes.
  • Training Losses. We use the sigmoid focal loss (for recognition) [60] and L1 regression loss (for localization) below. Existing works also use the same losses but only with fine labels; we apply the loss to both coarse and fine labels. Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes. Without our hierarchical loss, $C$=18. With our hierarchical loss, $C$=22 (18 fine grained + 3 coarse + 1 object class). $\lambda$ is set to 0.25. Modifications for other detectors similarly follow.
  • Optimization. We train all LiDAR-only detectors for 20 epoch using an AdamW optimizer and a cyclic learning rate. We adopt a basic set of data augmentations, including global 3D tranformations, flip in BEV, and point shuffling during training. We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU. The training noise (from random seed and system scheduling) is $<$ 1% of the accuracy (standard deviation normalized by the mean).
  • Post-processing. We use non-maximum suppression (NMS) on detections within each class to suppress lower-scoring detections. In contrast, existing works apply NMS on all detections across classes, i.e., suppressing detections overlapping other classes’ detections (e.g., a pedestrian detection can suppress other pedestrian and traffic-cone detections).
Figure 12: We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO. In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance. Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
B More Visualizations
We present additional visualizations of our multi-modal late-fusion approach in Fig. 12. Our method correctly classifies geometrically similar but semantically different categories like adult versus stroller, bicycle versus personal mobility, child versus adult, and adult versus construction worker. We provide a video of our results in the supplement. Our project page contains a video demo.
C Analysis on Transformer 3D Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [9]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [35] and CMT-L [10] in Table VI.
First, we note that BEVFusion-L performs 2.1 mAP better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%. However, BEVFusion-L performs 4.6 mAP worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging. Surprisingly, CMT-L performs worse than CenterPoint on all classes. We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2 mAP improvement overall, and a 8.5 mAP improvement for rare classes. This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy. In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3 mAP overall, and performs 5.2 mAP worse than our late-fusion approach. Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9 mAP improvement overall, with a 19.1 mAP improvement for rare classes. Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
@@ C Analysis on Transformer 3D Detectors
TABLE VI: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
@@
MMLF(CenterPoint, DINO) 51.4 77.9 59.4 20.0
@@
TABLE VII: Comparison of inference time. Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from the main paper for reference. Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time!
FCOS3D BEVFormer CP TransFusion DeepInteraction MMLF(CP, DINO)
[13] [23] [12] [34] [59] (Ours)
Time (ms) 89 327 323 367 590 323
mAP 20.9 27.3 39.2 39.8 43.7 51.4
D Inference Runtime
Table VII compares the inference runtime of our method against prior works on a single GPU. Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [23], TransFusion [34] and DeepInteraction [59]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including CenterPoint + DINO and [9] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
TABLE VIII: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and PO is police-officer. We highlight classes with Medium and Few examples per class in blue.
CenterPoint Multi-Head Car Ped. Barrier TC Truck Bus Trailer CV MC Bicyc.
Original 87.7 87.7 70.7 74.0 63.6 72.7 45.1 26.3 64.7 47.9
89.1 88.4 70.8 74.3 64.8 72.9 42.0 25.7 65.9 53.6
for LT3D 82.4 62.0 60.1 49.4 55.7 28.9 19.7 48.9 33.6
88.1 72.4 72.7 62.7 70.8 40.2 24.5 62.8 48.5
Adult PP CW Debris Child Stroller PO EV PM All
Original 64.0
64.8
for LT3D 81.2 21.7 14.2 1.1 0.1 0.1 1.3 0.1 0.1 31.2
86.3 32.7 22.2 4.3 0.1 4.3 1.8 10.3 0.1 39.2
E Analysis on Class Grouping
Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing. For example, CenterPoint groups pedestrian and traffic-cone since these objects are both tall and skinny. We study the impact of grouping for both the standard and LT3D problem setups. We define the groups used for this study below. Each group is enclosed in curly braces. Our group-free head includes all classes into a single group.
  • Original: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Pedestrian, Traffic Cone}
  • LT3D: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Adult, Child, Construction Worker, Police Officer, Traffic Cone}, {Pushable Pullable, Debris, Stroller, Personal Mobility, Emergency Vehicle}
We use the class groups proposed by prior works [18, 12] for the standard benchmark and adapt this grouping for LT3D. Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. The group-free approach simplifies architecture design, while also providing competitive performance.
Two insights allow us to train the group-free architecture. First, we make the group-free head proportionally larger to train more classes. The standard grouping setup contains 6 heads, each with 64 convolutional filters. Scaling up to the nearest power of two, our group-free head has 512 convolutional filters. Second, we do not perform between-class NMS. The standard setup performs NMS between classes in each group (e.g., since pedestrians and traffic cones are tall and skinny, the model should only predict that an object is either a traffic cone or a pedestrian). However, performing NMS between classes requires that confidence scores are calibrated, which is not the case. Moreover, for LT3D, score calibration becomes more important for rare classes as these classes have lower confidence scores than common classes on average, meaning that common objects will likely suppress rare objects within the same group. Our solution is to only perform within-class NMS, which is standard for 2D detectors [86].
TABLE IX: Different variants achieve similar performance. We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy. Unlike [87, 88] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
Method Hier. Many Medium Few All
CenterPoint (w/o Hier.) [12] n/a 76.4 43.1 3.5 39.2
CenterPoint w/ Hier. (a) 77.1 45.1 4.3 40.4
(b) 76.4 45.0 5.3 40.5
(c) 76.5 45.2 5.2 40.6
(d) 74.5 43.5 5.6 39.5
F Analysis of Hierarchical Training
Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [88]. We implemented such an approach, but found the training did not converge. Interestingly, [88] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted. Instead, we found better results using the method from [87] (a winning 2D object detection system on the LVIS [52] benchmark) which multiples class probabilities of predictions (e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax. We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference). Table IX compares their performance for LT3D.